### v2

# AI sustains higher strategic tension than humans in chess

Abstract
Strategic decision-making requires balancing immediate opportunities against long-term objectives: a tension fundamental to competitive environments. We investigate this trade-off in chess by analyzing the dynamics of human and AI gameplay through a network-based metric that quantifies piece-to-piece interactions. Our analysis reveals that elite AI players sustain substantially higher levels of strategic tension for longer durations than top human grandmasters. We find that cumulative tension scales with algorithmic complexity in AI systems and increases linearly with skill level (Elo rating) in human play. Longer time controls are associated with higher tension in human games, reflecting the additional strategic complexity players can manage with more thinking time. The temporal profiles reveal contrasting approaches: highly competitive AI systems tolerate densely interconnected positions that balance offensive and defensive tactics over extended periods, while human players systematically limit tension and game complexity. These differences have broader implications for understanding how artificial and biological systems navigate complex strategic environments and for the deployment of AI in high-stakes competitive scenarios.
Chess — Artificial Intelligence — Complex Networks — Strategic Tension
preprint: APS/123-QED
The aphorism that one may have won the battle but lost the war is encapsulated in the notion of a “Pyrrhic victory.” Costly short-term gains must be balanced against the longer-term uncertainties, opportunities, or challenges that may emerge in competitive environments. This presents a fundamental challenge in strategic decision-making for both biological and artificial agents across diverse competitive and cooperative landscapes [42, 40, 3], of which war [weinsteinRebellionPolitics2007, bohorquezCommonEcology2009] and conflicts [smithLogicAnimal1973, dedeoInductiveGame2010a, brushConflictsInterest2018] are poignant examples. In the more general case, organisms and individuals learn such as during development (e.g., neural pruning [woudeNoGains2019]), foraging (e.g., depending on climate [piteraDailyForaging2018]), and social change (e.g., reorganization of social hierarchy [28]) or, over evolutionary time scales, “adapt” to changing ecological niches [24]. In the context of an individual’s lifetime, these can be characterized as strategic choices that involve an implicit trade-off between overfitting the local environment and maintaining the flexibility to adapt to unforeseen environmental conditions [27], in an echo of the speed-accuracy trade-off [33]. In many real-world settings, a complete measurement of the strategic environment, costs, and outcomes is not available. Chess, by contrast, is a game of perfect information, providing a well-studied and bounded setting.
As a result, chess has served as a model system for investigating strategy. It has well-defined rules, perfect information, complexity, and rich history as a testbed for both human expertise and artificial intelligence (AI) [36, 39, 11, 32]. Chess is a competitive game between two players, white and black, who take turns making moves with a single piece on a board with 8x8 squares as shown in Fig. 1. Both players start with 16 pieces, and the objective of the game is to capture the enemy’s king. A key feature that makes the game interesting is the combinatorial explosion in potential moves, which prohibits exhaustive calculation. As a result, human players must make heuristic decisions about the possible long-term value of moves [39]. The nature of the trade-off is shaped by cognitive limitations, psychological pressures, and time constraints, often leading to biased or suboptimal decisions [21, 25, 20]. In contrast, modern AI exhibits superhuman performance through two main strategies. One class involves deep reinforcement learning systems, such as AlphaZero and Leela Chess Zero, which develop strategies through self-play using deep convolutional neural networks. Alongside these, engines like Stockfish combine massive brute-force search trees with NNUE (Efficiently Updatable Neural Networks). While Stockfish originally relied on handcrafted evaluation functions, modern versions utilize machine learning to train their evaluation layers, blending classical search efficiency with neural network-based intuition. These differences suggest that the nature of the trade-offs that AI engines make is distinct from those made by human players. Still, there is little characterization of the properties that distinguish gameplay.
Here, we compare humans to AI play by analyzing the network structure of piece-to-piece interactions [16] over many chess games. We consider thousands of games across a broad range of scenarios, including games with players of all skill levels and official AI tournaments between competitive chess engines, such as Leela Chess Zero and Stockfish, held between 2019 and 2024. In these tournaments, the engines operate adaptively, allocating varying amounts of time to different moves. In addition to these tournament games, we include simulations in which Stockfish is run at fixed search depths. This serves as a controlled baseline, allowing us to evaluate piece-piece interaction networks under uniform computational constraints.
We find that the complexity of the game, as measured by “strategic tension,” follows, on average, a regular curve, whose features correspond to the different stages of the game. The curve is characteristically different for human players compared with AI players. Our observations suggest that humans and AI exhibit distinct behaviors in complex strategic environments, raising questions about the consequences of delegating competitive real-life scenarios to AI.
I Defining Tension in Chess
In chess, tension refers to a state of unresolved tactical interaction where multiple pieces from both sides occupy squares from which they can capture or be captured by one another. Unlike a direct exchange of pieces, which resolves the interaction and simplifies the board state, tension represents a “latent” complexity. It persists as long as players maintain these mutual threats without acting upon them, forcing both competitors to constantly evaluate the shifting consequences of a potential explosion of captures. By formalizing these threats into a network of offensive and defensive relationships, we can quantify this tension as the density and structure of piece-to-piece interactions.
Given the state of a chessboard as in Fig. 1a, we consider the set of offensive and defensive relationships at each ply (half-move). An attacking piece can capture an opposing piece in a single move. The defending piece would threaten any piece that would capture the piece it is defending. This is the corresponding weighted network in Fig. 1b, where a node is either a chess piece $p_{i}$ or a vacant square $s_{i}$, and the links can be of three types:
  • Attack link: We define an attack link from $p_{i}$ to $p_{j}$ if $p_{i}$, a piece from one side, can legally capture $p_{j}$, a piece from the opposing side. Attack links reflect direct threats, highlighting where immediate exchanges could occur.
  • Defense link: We define a defense link from $p_{i}$ to $p_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally capture $p_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ in a single move. Kings can defend other pieces, but cannot be defended under this definition. This distinction arises because the King is never captured in chess; hence, the legal mechanism of a reciprocal capture (or “re-capture”) cannot apply to it. Defense links reveal the layers of protection that influence whether an attack can be safely carried out or needs to be postponed.
  • Control link: We define a control link from $p_{i}$ to a vacant square $s_{j}$ if there exists an opponent’s piece $p_{k}$ that can legally move to $s_{j}$ and after $p_{k}$’s move, $p_{i}$ would be able to legally capture $p_{k}$ on $s_{j}$ in a single move. Control links, finally, represent subtler forms of influence, where empty squares become contested zones.
In constructing the network, we disregard which side is to move, focusing instead on the full set of interactions at each ply to capture the full state of the board. Attack and defense links can have weight 1 or 2, depending on whether the corresponding conditions are satisfied by one or both of the pieces involved. In contrast, control links always have weight 1. By construction, the network has at most 64 connected nodes, each corresponding to either a chess piece or a vacant square. For our work, we do not distinguish between link types  [^1] Although we also experimented with alternative networks that distinguish between link types and assign values according to the pieces involved, following commonly accepted conventions, these variations yielded qualitatively similar results in our case studies. We therefore opted for the simpler representation to maintain interpretability and focus on the main structural features of the positions.. The resulting network captures the interplay among pieces and squares that creates and sustains tension on the chessboard.
Figure 1: One example of tension network. (a) Chessboard at the ply 44 during a game between the two grandmasters, Magnus Carlsen (white) and Alexey Sarana (black), in the Titled Tuesday tournament, Round 6, held on chess.com on November 26, 2024. (b) The network represents the interactions on the chessboard. The chess pieces are depicted as dark and light circles, depending on whether they are black or white. Attack links are displayed in orange when originating from white pieces and in red when originating from black. Defense links are turquoise when white is defending and blue otherwise. Control links are light blue (originating white) and light green (originating black). Small, brownish nodes represent vacant squares under control. The chosen ply represents the highest tension reached during the game according to the maximum eigenvalue of the adjacency matrix.
Among the many ways to characterize the interaction network, we focus on how game complexity evolves, reflected in features such as the number of interacting pieces and loops  [^2] A loop consists of a closed sequence of interactions between pieces forming a cycle within the network (e.g., a white rook on b4 defends a white knight on e4; the knight, in turn, attacks a black bishop on d6, which attacks the rook on b4).. We interpret this complexity both in terms of structural richness and strategic fragility, namely the potential for cascading consequences following a suboptimal move. In dense, highly connected configurations, even a small error may propagate across the network, shifting advantage substantially. To capture this latent vulnerability, we use the maximum eigenvalue of the adjacency matrix, $\lambda_{1}$, adopting the chess-inspired term strategic tension. This quantity summarizes how tightly coupled the interaction network is at any moment and reflects critical structural dynamics [31]. For simplicity, we consider the undirected tension network, treating tension as a reciprocal property between pieces.
The maximum eigenvalue reflects network functionality, encompassing both its structural features and dynamical properties [8]. This can be understood via topological entropy, which captures the complexity and unpredictability of paths in a network [1]: the spectral radius summarizes the network’s “diversity” of possible trajectories. Higher $\lambda_{1}$ indicates greater dynamical complexity, consistent with tension as a measure of uncertainty or potential instability. Importantly, $\lambda_{1}$ is bounded between the average and maximum node degree, $\langle s\rangle\leq\lambda_{1}\leq s_{\max}$, linking global connectivity with local heterogeneity [41]. Thus, the spectral radius serves as a natural metric for tension, bridging structural properties (as ensured by the Perron–Frobenius theorem) and dynamical behavior (as reflected in entropy-like measures) at the current state of the board (see the Supplementary Materials for details).
RESULT 1: humans peak, AI sustains STRATEGIC TENSION
We illustrate the typical evolution of the strategic tension across multiple games in Fig. 2. We can distinguish four main phases. The first phase, corresponding roughly to the opening stage up to around ply 30, features a buildup where pieces approach each other, and contested squares come under threat, leading to an abrupt increase in tension. The second phase, spanning approximately from ply 30 to ply 60, represents peak complexity, where the board remains highly interconnected, and the tension stays elevated. The third phase exhibits a gradual simplification of the tension network, as piece exchange and loss steadily lower the overall tension. Finally, the games reach a relatively stable value of tension, but at this point available statistics become increasingly sparse, especially for games played between humans that typically end earlier (Fig. 2b and 2c). Thus, the dynamics of tension capture the overall contours of the game, including the opening buildup, the combinatorially complex midgame, and the relaxation leading up to the endgame as discussed in the theory of chess play [13].
Despite the overall similarities in the typical trajectory of strategic tension between human and AI play, there are notable differences. Maximum tension is higher in human-played games than in engine-played games. Up to approximately ply 40, human players typically follow established opening repertoires from theoretical databases, which often lead to well-studied positions with numerous direct piece interactions, including attacks, defenses, and immediate threats [15, 6]. These theory-driven positions generate the highest points of tension as players transition to original play in the midgame. Once players leave the realm of memorized theory and must rely on their own calculation and evaluation, they tend to simplify positions. Human players reduce tension more rapidly and, as we discuss below, resolve piece conflicts sooner, often favoring attack-oriented configurations. In contrast, state-of-the-art engines tend to maintain more distributed, positional advantages that unfold gradually over multiple moves, resulting in a slower decay of tension.
A secondary difference emerges in how strategic tension declines in drawn vs. decisive games. Drawn games between humans show a rapid and substantial fall in tension relative to decisive games. This agrees with the faster decrease in the number of pieces for drawn games between human players (see Fig. 10 in the Supplementary Materials). AI games show a contrasting pattern. Drawn games show significantly more (not fewer) pieces on the board relative to decisive games in the endgame and tension does not differentiate strongly.
II Result 2: Strategic tension reflects, does not predict game result
To relate the structural complexity of the interaction network to game outcomes, we compare the evolution of strategic tension with the algorithmic evaluation provided by Stockfish (see Fig. 8 in the Supplementary Materials). Inspecting the evaluation score, we identify bifurcation points in game trajectories, typically around ply 30 (move 15), where decisive and drawn games begin to diverge for human players during the opening-to-midgame transition. A similar behavior is observed in AI games, where a local maximum or inflection point in the evaluation profile precedes the divergence of outcomes. Notably, in all games, this divergence occurs before strategic tension reaches its absolute peak. This ordering suggests that the game’s outcome is largely determined during the ascent toward peak tension, before the subsequent reduction in tension differentiates decisive games from draws. In chess terms, foundational advantages—such as positional pressure, space, or superior piece coordination—are established while the board is still increasing in complexity.
The ensuing reduction in tension reflects a phase of conversion or simplification from the tension peak, often marking the resolution of a struggle in which one side has attempted to generate complications. Consistent with established principles of high-level play, players with a decisive advantage tend to trade pieces to limit counterplay [37]. Our results confirm that this reduction is not the cause of the outcome, but rather its consequence: both humans and engines simplify positions once a win or a forced draw is already favored.
Figure 2: Tension values during games for both humans (grandmasters) and AI (Stockfish and Leela Chess Zero). We consider 1,200 games for each, with human games obtained from https://www.pgnmentor.com and AI games from https://tcec-chess.com, which hosts official engine-versus-engine tournaments (a) Average tension is represented using cold and warm colors for humans and AI, respectively. For both groups, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively. (b) For human games, the plots show the percentage of games—relative to the original total—that end in a draw or in a decisive result (win/loss) at each ply. Initially, the dataset is split between draws and decisive outcomes, and both percentages steadily decline as the number of plies increases, indicating how long games tend to last statistically. (c) The same analysis is performed for AI games, revealing a similar trend.
RESULT 3: Variability in strategic tension reflects different gameplay
While tension seemingly captures only general and coarse features of the game, it is correlated with key aspects of gameplay that distinguish humans from AI (see Fig. 9 in the Supplementary Materials).
Number of links and loops: AI interaction networks exhibit more links and loops, indicating a richer web of mutual threats, defenses, and controls. However, the maximum average number of loops is higher in human games, reflecting the peak of human tension.
Number of pieces: AI maintains a higher number of pieces over the trajectory of games on average. This reflects a weaker tendency to simplify the game by removing pieces from the board. From mid to endgame, the small number of pieces limits the strategic tension that is typically reached in human games.
Figure 3: Comparison of structural properties across human and AI games. We analyze the set of games shown in Fig. 2. Two measures are considered: the average standard deviation of the degrees associated with chess pieces, normalized by the mean of the degrees, and the average tension per piece. Red is for AI and blue for human games.
Degree distribution: The noise coefficient of the degree distribution over chess pieces (defined as the standard deviation over the mean degree) mirrors the rise and fall of strategic tension, first falling then rising. While human and AI games display a broadly similar global trend, human play exhibits more pronounced variation as shown in Fig. 3a. In particular, during the opening phase, the relative degree variation decreases markedly, reflecting a positioning phase in which interactions are more evenly distributed across pieces as tension accumulates. As the game progresses, however, the interaction network in human games exhibits greater variability, eventually returning to levels similar to those observed in AI games, or even lower.
Tension per piece: As shown in Fig. 3b, the tension per piece $\lambda_{1}/n_{t}$ remains consistently higher across all plies in human games. While human play shows an approximately linear increase in tension per piece throughout the game, AI games display a sublinear progression that appears to plateau around ply 80. Human games involve more frequent captures of pieces (see Supplementary Materials), but the remaining pieces are then involved in a greater number and greater complexity of interactions.
Attack-defense balance: When distinguishing between attack and defense links, we find that the surplus of attack links over defense links, or “attack-defense balance” as shown in Fig. 4, moves inversely with the tension profile. It first drops in the early phase more defense-oriented phase of the game. As the game progresses, attack links gradually outnumber defense links. The curve is shifted towards more positive values in human games compared to AI games.
Figure 4: Balance between attack and defense links across human and AI games. We analyze the set of games shown in Fig. 2. For both human and AI games, we distinguish between games that end in a draw and those that result in a win/loss, represented with light and dark shades, respectively.
Overall, we find that strategic tension aligns with basic gameplay metrics, such as the number of pieces and, in particular, the presence of loops in the interaction graph. At the same time, it is informative to relate it to the balance across pieces on the board and to how pieces are arranged in formations. Taken together, our results indicate that AI games retain more pieces, exhibit a more balanced network of attack–defense connections, and display less variability in the degree distribution across pieces.
RESULT 4: Cumulative TENSION SCALES WITH SKILL AND COMPUTATIONAL COMPLEXITY
Figure 5: Tension loads for different Elo ratings and Stockfish depths. (a) We represent tension loads (sum of tension values up to ply 150) for rapid games (turquoise dots), blitz games (green dots), classical games played by top human players (blue star), and games between Stockfish and Leela from official TCEC tournaments (red star), all associated with estimated Elo ratings. Each dot/star represents the average across 1,200 games. The blue and red stars were derived from the data analyzed in Fig. 2. For reference, blitz games are played with very short time controls (typically 3–5 minutes per player), rapid games with intermediate time controls (around 10–25 minutes per player), and classical games with long time controls (often 60 minutes or more per player) allowing deeper strategic play. (b) We represent tension loads from games between versions of Stockfish at varying depth levels (orange dots), computed from a smaller sample of 120 games per dot. The two panels also show the $R^{2}$ values of the corresponding linear regressions.
While engines can maintain higher tension for a longer duration compared to humans, does this difference depend on skill level? We consider the cumulative tension up to ply 150 across different Elo ratings for human players in Fig. 5a  [^3] In principle, Stockfish has a search depth parameter that serves as a proxy for skill level. Leela has a time limit parameter during which it is permitted to assess the board state and make a decision. At a few dozen search depths, the tension load is limited to about 450, similarly for Leela at a few seconds. Since the games that are shown go far beyond these limited parameter values with supercomputers, this suggests that algorithmic complexity increases the ability to handle tension.. Tension changes substantially with Elo, showing an approximately linear dependence in human games. Moreover, across all Elo values, rapid games consistently exhibit slightly higher tension loads than blitz games. This is consistent with the idea that longer time controls allow human players to create and manage more complex positions [5]. The effect is even more pronounced in classical games, where players have significantly more time to make decisions, resulting in higher tension and reflecting the chess community’s concern that time-constrained players experience larger performance drops on complex decisions and therefore tend to choose safer, less strategically risky continuations [5, 7]. In contrast, elite classical play is often characterized by keeping simplification at bay and maintaining a complex struggle, where grandmasters sustain tense, fluid positions that test the opponent’s stamina and depth of calculation over many hours [22, 30, 12].
For comparison, we examine how the tension load changes for Stockfish, where we can control the computational complexity of the algorithm by adjusting the search depth at a fixed value for simulated games in Fig. 5b. Although we examined search depths only up to 50, the results already indicate that increasing depth leads to higher tension. As a reference, previous experiments using a different engine (Houdini) reported a roughly linear relationship between Elo rating and search depth up to 20 plies [17]. Overall, this indicates that increased tension load is indicative of greater computational demands on the player.
III Discussion
We investigate the dynamics of strategic tension, a network-based definition of piece interactions, in chess games played by humans and AI. Our central hypothesis is that distinct tension profiles would emerge, reflecting differing strategic approaches. Our observations confirm this hypothesis and reveal that state-of-the-art chess engines (Stockfish at high search depths and Leela Chess Zero at advanced neural network settings) consistently maintain significantly higher levels of strategic tension for longer durations than even the best human grandmasters.
We speculate that the sustained tension in AI play reflects the deferral of immediate opportunities and facilitates latent strategic options that only materialize in later phases of the game. This kind of strategic decision making would be facilitated by the ability to handle greater computational complexity. Correspondingly, we find that human players exhibit a linear relationship between cumulative tension and skill level (Elo rating). Similarly, Stockfish’s tension load increases with computational complexity, as reflected in its search depth. Far beyond these examples are AI engines running on supercomputers. Superhuman computational capacity obviates the need to simplify upon leaving theoretical territory; AI engines do not need to rely on heuristics and intuition such as familiarity with positions in opening theory that heavily guide expert human play. These observations suggest that the ability to sustain strategic tension is a hallmark of (for human players) advanced chess expertise, encompassing deeper strategic understanding, and extensive theoretical knowledge and (for AI) enhanced calculation skills [12, 9].
The increase in tension with computational load likely arises from different factors in the two groups, reflecting their nature as artificial versus biological systems. In AI, tension rises as engines gain computational capacity and play at higher levels, where controlling more pieces and squares maximizes options and exploits opponents’ mistakes [4]. In humans, more skilled players can handle complex positions and maintain higher‑tension situations without simplifying [23, 11, 10, 18]. Indeed, humans may tend to simplify under certain conditions possibly due to limitations in working memory, attention, or susceptibility to psychological pressure [20, 26]. This interpretation of the results could also suggest that human players may sometimes actively steer games toward drawn resolutions, reflecting deliberate risk minimization and metagame considerations typically absent in AI play. The simplification of high-tension positions thus serves not only strategic purposes but also as an adaptive response to the computational demands of accurate tactical calculation, allowing players to sustain the level of accuracy needed for strong play within cognitive limits [14]. Of course, such limits come from the evolutionary trade-offs reflected in the brain, where short-term costs, such as losing a chess game, are outweighed by long-term benefits in metabolic efficiency and general, rather than domain-specific, intelligence [snell-roodBrainSizeGlobal2009, burgerAllometryBrainSize2019].
Several avenues for future research emerge from this work. Extending this methodology to other complex strategic games, such as Go [38], Shogi, or even real-time strategy e-sports, could test the universality of the observed dynamics across different game structures (e.g., imperfect information and varying spatial dynamics). Moreover, the AI engine used (Stockfish) represents one architecture; other engines with different algorithms, such as neural network-based engines like AlphaZero, may show different tension dynamics. While we focused on Stockfish and Leela due to their prominence and availability, our analytical framework and open-source code can readily be applied to other UCI-compatible engines, enabling broader comparative studies. With a more direct focus on tension, directly investigating the cognitive and physiological correlates of our tension metric in human players—for example, by combining game analysis with eye-tracking, think-aloud protocols, or stress indicators such as galvanic skin response—could help elucidate the underlying mechanisms driving changes in tension.
The observed differences between artificial and human systems could have implications for AI designed to interact with humans, particularly as AI is expected to play an increasing role in teaching and training across many domains. In chess, a notable step in this direction is the Maia engine [29], which is explicitly designed to model human play rather than engine-optimal moves. Moreover, while direct extrapolation requires caution, examining how tension is sustained in other complex adaptive systems, such as economic markets or geopolitical negotiations, can provide insight [35]. The finding that increased computational complexity leads AI to sustain tension may matter for stability in international relations [simmons-edlerPositionAIPowered2024], potentially enabling more complex standoffs but also increasing the stakes of miscalculation [19, guttieriIntegrativeComplexity1995, 34], creating an AI-enabled parallel to Dr. Strangelove.
Code availability
Python code used to compute tension and generate the Stockfish self-play games is available on GitHub at https://github.com/Draghellon/Tension_Chess.
Authors’ contributions
AC led the definition of the tension metric, with contributions from the other authors, and carried out the code development, numerical analyses, and figure generation. EDL performed several simulations on a computing cluster using Stockfish at high depths. VDPS provided the original inspiration for the project, supervised its development at all stages, and independently cross-checked the results. All authors participated in the discussions and contributed to writing the manuscript.
Acknowledgements
We thank V. Latora for valuable discussions, particularly regarding the inclusion of Fig. 4, and M. Barthelemy for useful discussions and for sharing his code, which we used as a baseline for our analysis. This work was partly funded by the City of Vienna, Municipal Department 7, and the Federal Ministry of the Republic of Austria for Climate Action, Environment, Energy, Mobility, Innovation, and Technology as part of the project GZ2021-0.664.668. It was also partially funded by the Austrian Science Fund (FWF) 10.55776/ESP127.
IV Methods
IV.1 Data
Chess games are commonly stored in the PGN (Portable Game Notation) format. This standard plain-text representation records the sequence of moves in a game, along with optional metadata such as player names, event, result, and time controls. This format is human-readable and machine-parsable, making it ideal for large-scale game analysis.
The games between top-level human players were obtained from https://www.pgnmentor.com, which provides thousands of PGN-formatted games organized by player or sorted by opening. This source includes a comprehensive archive of historical and modern games, making it well-suited for comparative analysis.
Competitive matches between Stockfish and Leela Chess Zero were collected from the https://tcec-chess.com/ website, which regularly publishes high-level engine-versus-engine games in PGN format. These games represent the cutting edge of chess engine performance under standardized tournament conditions. The typical time control is 120 minutes per player, with a 12-second increment added after each move.
These datasets served as the basis for the analyses presented in Fig. 2.
In addition, a large set of rapid and blitz games played by human players with Elo ratings ranging from 1000 to 2700 was downloaded from https://lichess.org/, an online platform that provides a vast archive of games. This dataset broadly represents amateur to expert players and is ideal for analysis because games are subject to anti-cheating controls, ensuring reliable data across a broad skill spectrum.
Together with the previously described datasets, these data formed the basis for the analyses presented in Fig. 5.
IV.2 Software
All operations related to chess move generation, rule validation, and position analysis were carried out using the python-chess library (v1.9.4). This comprehensive library offers tools for representing chess boards, parsing PGN files, and computing legal moves according to FIDE rules. Furthermore, it provides native support for communicating with UCI-compatible engines such as Stockfish, which allowed us to automate the evaluation of positions and simulate engine-vs-engine games with varying search depths.
IV.3 Generating Stockfish games at varying depths
To generate the games between Stockfish and itself at various search depths, we followed a procedure designed to balance realism and controlled variability. While there is no universally accepted mapping between Stockfish’s search depth and Elo rating, approximate correspondences have been estimated in prior works [17]. Based on these estimates, we selected a range of depths roughly aligned with the playing strength of top human players and beyond.
To introduce diversity in the opening phase and avoid deterministic repetition, each game was initialized using the first 12 plies (typically covering the opening) from randomly selected games played by top-level human players. After this opening phase, the rest of the game was played entirely by two instances of Stockfish, one playing each move at either depth $D$ or $D+1$, chosen randomly, and the other at either depth $D$ or D1D-1, also chosen randomly. This asymmetry introduced slight variability in evaluation and decision-making while preserving a consistent overall depth level $D$.